OpenStudy (konradzuse):

4.10 #1. In Exercise 1 let T[a] and T[b] be the operators whose standard matrices are given. Find the standard matrices for T[a] o T[b] and T[b] o T[a]

OpenStudy (konradzuse):

A =\[\left[\begin{matrix}1 & -2 & 0 \\ 4 & 1 & -3 \\ 5 & 2 & 4\end{matrix}\right]\] B = \[\left[\begin{matrix}2 & -3 & 3 \\ 5 & 0 & 1 \\ 6 & 1 & 7\end{matrix}\right]\] @satellite73

OpenStudy (unklerhaukus):

what do you mean by the little circle o

OpenStudy (konradzuse):

it's just what the book says, let me screen shot.

OpenStudy (unklerhaukus):

what does it mean

OpenStudy (konradzuse):

I believe it says the composite of linear transformations.

OpenStudy (konradzuse):

http://tutorial.math.lamar.edu/Classes/LinAlg/LinearTransformations.aspx This has it at the bottom as well, it's not really explaining how we get to an answer... I see from the previous section 4.9 that we use rotations and such but.....

OpenStudy (konradzuse):

Save me :).

OpenStudy (konradzuse):

It gives me the answer already, but I want to know why that's the answer and how to get it...

OpenStudy (unklerhaukus):

so \[T_A \circ T_b = T_A ( T_b(x) )\]

OpenStudy (konradzuse):

mhm

OpenStudy (konradzuse):

but there is no x here?

OpenStudy (konradzuse):

nor does it say what kind of transformation we are performing...?

OpenStudy (unklerhaukus):

try this \[T_A\circ T_B=\text A\text B=\left[\begin{matrix}1 & -2 & 0 \\ 4 & 1 & -3 \\ 5 & 2 & 4\end{matrix}\right]\left[\begin{matrix}2 & -3 & 3 \\ 5 & 0 & 1 \\ 6 & 1 & 7\end{matrix}\right]=\]

OpenStudy (konradzuse):

\[\left[\begin{matrix}-8 & -3 & 1 \\ -5 & -15 & -8 \\ 44 & -11 & 45\end{matrix}\right]\]

OpenStudy (konradzuse):

wow... Is that all we needed to do? That's some BS LOL.....

OpenStudy (unklerhaukus):

looks good

OpenStudy (konradzuse):

Why is that all we have to do... Why was that so easily konfusing... :'(. I hate Linear >(